OpenStudy (anonymous):

PLEASE HELP! TEST TOMORROW! tell whether the triangle is right, obtuse, or acute 5n, 12n, 13n where n >0

Parth (parthkohli):

Does this follow the Pythagorean Theorem?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

im in geometry but yeah its the pythagorean theorem

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

what kinds of triangles follow the pythagorean theorem?

OpenStudy (anonymous):

right, obtuse, and acute

Parth (parthkohli):

hmm, not quite.

OpenStudy (anonymous):

yes. its a^2+b^2=c^2 and its asking to tell if the triangle is right, obtuse, or acute

OpenStudy (anonymous):

@mathmale please help!!

Parth (parthkohli):

only a special kind of triangles follow the pythagorean theorem.

OpenStudy (anonymous):

ok well how do you do the problem then

OpenStudy (anonymous):

5n^2+12n^2=13n^2 doesnt give the right answer

OpenStudy (anonymous):

@ParthKohli @mathmale

OpenStudy (anonymous):

@ParthKohli can you please tell me how to do this

Parth (parthkohli):

that's not 5n^2 + 12n^2 = 13n^2. that's (5n)^2 + (12n)^2 = (13n)^2

OpenStudy (anonymous):

got it thanks

Parth (parthkohli):

you'll see that both sides are the same. so the pythagorean theorem is followed. pythagorean theorem is true for right triangles only.

OpenStudy (anonymous):

yeah this one is right thanks